# arXiv:2602.23010v2[cs.GR] 01 Mar 2026

### v2

@@ 6.2 Cross-Validation
On He et al. 2022 (regularizer, 82 pairs), Helmlab achieves stress 30.3 vs CIEDE2000’s 32.6. On MacAdam 1974 (held out entirely, 128 pairs), Helmlab achieves 20.3 vs CAM16-UCS’s 18.7 (best). Figure˜4 visualizes these results. Helmlab outperforms CIEDE2000 on He et al. but trails CAM16-UCS on MacAdam. This is consistent with what one would expect from a model trained primarily on combvd: it generalizes reasonably but does not dominate on all evaluation sets.
@@ 6.3 Generation Properties: The Measurement–Generation Tradeoff
A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation. During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.34$. The optimizer had deformed the achromatic axis to reduce stress, which made gray gradients unusable because they showed visible color artifacts. Adding a penalty to constrain grays to $C=0$ did not help: the optimizer either stayed in the measurement-optimal basin or escaped to a different basin with much worse stress (26–31). There was no smooth transition between the two.
The neutral correction (Section˜3, Stage 10) addresses this problem. It subtracts the achromatic error after all enrichment stages, so the optimizer can find its measurement-optimal parameters without constraint. The achromatic axis is then corrected separately. The stress increase from the neutral correction alone is +0.04+0.04 (from a measurement-only baseline of 23.18 to 23.22); the subsequent blue-band refit (Section˜6.6) adds another +0.08+0.08, yielding the final 23.30. Achromatic chroma drops from 0.34 to $10^{-6}$.
Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage. Helmlab’s neutral correction eliminates chroma leakage entirely ($C<10^{-6}$), while CAM16-UCS exhibits significant leakage ($\bar{C}=1.42$). Table˜2 summarizes all generation properties.
Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC guarantees $C<10^{-6}$ for all grays.
Table 2: Generation quality metrics.
@@
Achromatic max $C$ $10^{-6}$
Hue RMS (6 primaries) 18.1°
Hue max error 23.9°
Round-trip $\|\cdot\|_{\infty}$ $10^{-14}$
Munsell CV 20.0%
Jacobian det (min) 0.10
@@
5-fold CV gap (mean) +0.98
He stress 30.33
MacAdam stress 20.33
The achromatic guarantee means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 18.1° means sRGB primaries land within about 24° of their expected positions. This is acceptable for palette generation but not perfect. The rigid rotation (Section˜3, Stage 11) provides this hue improvement without affecting the distance metric because the metric is invariant under rotation. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
@@ 6.4 Ablation Study
Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their optimized values. This measures the marginal contribution of each component given the current parameter setting, not its isolated effect.
Table 3: Ablation: effect of removing components (frozen, not re-optimized).
Configuration stress $\Delta$
Full Helmlab 23.30
Euclidean distance only 27.60 +4.3
No H-K embedding ($w_{\mathrm{HK}}=0$) 27.30 +4.0
No hue correction 38.81 +15.5
No dark L compression 23.49 +0.2
Rotation (-28.2°) 23.30 0.00
No rotation 23.30 0.00
With Euclidean distance, the space alone gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full metric brings it down to 23.30, recovering a further 4.3 points. Compared to other Euclidean baselines (CAM16-UCS = 33.9), the space alone gives a substantial improvement. Among individual stages, hue correction is the most critical ($\Delta=+15.5$), followed by the H-K embedding ($\Delta=+4.0$). These large frozen-ablation deltas reflect the co-optimization of all parameters: disabling one stage disrupts the balance of the entire pipeline. The rotation row confirms that the metric is exactly invariant under rotation.
@@ 6.5 Sub-Dataset Performance
Table˜4 breaks down stress by combvd sub-dataset. Helmlab outperforms CIEDE2000 on 3 of 6 sub-datasets (BFD-P(D65), BFD-P(M), Witt). On the blue-heavy sub-datasets (LEEDS, RIT-DuPont), the blue-band refit reduces the gap substantially compared to an earlier version without the penalty: RIT-DuPont improves from 23.49 to 20.41, nearly matching CIEDE2000’s 19.51. BFD-P(C) is the only sub-dataset measured under illuminant C (rather than D65) and with a 2° observer, which may explain the remaining deficit.
Table 4: stress on individual combvd sub-datasets.
Sub-dataset NN Helmlab CIEDE2000
BFD-P(D65) 2,028 23.11 24.14
BFD-P(M) 548 22.22 35.09
BFD-P(C) 200 30.15 30.05
Witt 418 28.88 30.37
Leeds 307 20.92 19.46
RIT-DuPont 312 20.41 19.51
6.6 Blue-Band Analysis
During development, we observed that an initial version of Helmlab (without the blue-band penalty) exhibited severe gradient non-uniformity in the blue–cyan region. A red-to-blue gradient interpolated in Helmlab Lab coordinates showed a $51.4\times$ max/min step-size ratio when measured with CIEDE2000, compared to $1.3\times$ for the same gradient in CIE Lab. Frozen ablation traced the problem to the 4-harmonic hue correction, whose Fourier coefficients overshoot in the blue–magenta band ($h\approx 240$340340°). Disabling hue correction reduced the gradient ratio from $51.4\times$ to $4.9\times$ but increased combvd stress by +15.5+15.5 points, making this an unacceptable trade-off.
The solution was to add explicit sub-dataset penalties to the loss function (Equation˜17) and re-optimize all 72 parameters. With $\lambda_{b}=0.1$, the red-to-blue gradient ratio drops from $51.4\times$ to $5.8\times$ ($8.9\times$ improvement), while combvd stress increases by only +0.08+0.08 (from 23.22 to 23.30). stress improves on 5 of 6 sub-datasets (Table˜4), with the largest gain on RIT-DuPont (3.08-3.08 points). The only regression is BFD-P(D65) (+0.49+0.49), the largest sub-dataset, which accounts for the slight overall stress increase.
@@ 7 Practical Adoption
Figure˜6 shows the sRGB gamut boundary in Helmlab coordinates at three lightness levels. The boundaries are smooth and well-behaved. Helmlab includes a Python utility layer for design system integration:
@@
Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB or Display P3 color, preserving hue and lightness.
@@
Lightness ramps and hue rings at fixed chroma produce perceptually-spaced palettes. Semantic scales (Tailwind-style 50–950) are generated by interpolating L between 0.97 and 0.10 with chroma/hue preserved.
@@
When surround parameters are trained, the surround $S$ directly modulates the transform. Currently, a fallback soft L-inversion maps colors between light ($S=0.7$) and dark ($S=0.2$) modes with contrast enforcement.
@@
A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON.
@@ 8 Limitations and Future Work
    Model complexity. 72 parameters is considerably more than Oklab ($\sim$15 designed parameters). Each stage has a perceptual motivation, but the overall model is harder to interpret than simpler spaces.
@@
    Hue alignment. RMS hue error of 18.1° is a compromise. Some hue regions (cyan) have errors up to 23.9°. Further harmonic terms could improve this at the risk of overfitting.
@@
    Surround dependency. The surround parameter $S$ is architecturally present but not yet trained on viewing-condition data. Current dark/light adaptation uses heuristics.
@@
    Sub-dataset variation. Helmlab underperforms CIEDE2000 on 3 of 6 combvd sub-datasets (Table˜4). The blue-band refit improves 5 of 6 sub-datasets compared to the previous version; only BFD-P(C), measured under illuminant C with a 2° observer, remains notably worse (30.15 vs 30.05).
@@
    Not general-purpose. Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing). It is not designed for print, photography, or spectral applications.
@@
    Blue-band trade-off. The blue-band refit (Section˜6.6) reduces gradient non-uniformity by $8.9\times$ at a cost of +0.08+0.08 stress on combvd. The Munsell CV increases from 17.6% to 20.0%, hue RMS from 16.1° to 18.1°, and the He et al. cross-validation stress increases from 29.0 to 30.3.
@@
The open-source release includes a second, generation-optimized space (GenSpace) with shared $\gamma\!=\!\tfrac{1}{3}$ and no enrichment stages, which trades distance accuracy for faster, invertible gradient and palette generation. For future work, we plan to train the surround parameter on paired light/dark viewing data and to expand the training set with targeted human feedback using the bidirectional optimization framework that is already in the codebase.

### v3

@@ 6.2 Cross-Validation
Figure˜4 visualises cross-dataset performance on academic held-out sets:
  • He et al. 2022 (82 pairs, 3D-printed): MetricSpace 35.9 vs CIEDE2000 32.6. v21 trails on this set; the loss is discussed above.
  • MacAdam 1974 (128 pairs, fully held-out, never seen by optimiser): MetricSpace 19.51 vs CIEDE2000 22.13. v21 also leads OKLab (32.72), $J_{z}a_{z}b_{z}$ (24.14), CIE Lab (24.53), and DIN99 (23.31). On this set CAM16-UCS is the strongest appearance-model baseline: it scores 18.71, leading v21 by 0.80 stress — CAM16’s appearance-model formulation matches MacAdam’s experimental conditions (D65 illuminant, constant Y, controlled-viewing) better than COMBVD’s industrial samples. CIECAM02-UCS (19.21) and CIE94 (19.78) sit between these. v21 leads the rest of the field by 3+ stress points on MacAdam while ceding the top spot to CAM16; over the three-dataset average (Table˜1, Avg) it remains stress 21.75, vs CIECAM02-UCS 35.98 and CAM16-UCS 36.73.
The 3-dataset sub-balanced loss (Equation˜16) widened the optimum into a region that generalises better to small held-out sets without the v20b-specific blue-band penalty, and it pulled the model toward the BFD-class large datasets at the cost of small-sample sets like He 2022.
@@ 6.3 Generation Properties: The Measurement–Generation Tradeoff
A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation. During development, the measurement-optimal configuration mapped neutral grays to chroma $\bar{C}\approx 0.29$. The optimizer had deformed the achromatic axis to reduce stress, which made gray gradients unusable because they showed visible color artifacts. Adding a penalty to constrain grays to $C=0$ did not help: the optimizer either stayed in the measurement-optimal basin or escaped to a different basin with much worse stress (26–31). There was no smooth transition between the two.
The neutral correction (Section˜3, Stage 10) addresses this problem with a deliberate two-mode toggle. The correction subtracts the achromatic error after all enrichment stages, which allows the optimiser to find its measurement-optimal parameters without any achromatic constraint. The trained checkpoint is then used in two ways:
  • Distance-prediction mode (NC off). The headline stress of 22.48 is reported with the neutral correction disabled: the LUT subtraction is non-uniform along the chromatic plane and pulls some non-grey pairs into a slightly different distance basin, costing $\sim$+6.2+6.2 stress on combvd. We therefore evaluate the distance metric against the trained-basin coordinates directly.
  • Generation/authoring mode (NC on). For every UI workflow that touches an achromatic value (gray ramps, palette interpolation, dark/light token derivation, contrast checking on neutral foregrounds), the neutral correction is enabled and the maximum chroma for grays drops from $\sim 0.29$ to $10^{-5}$ on the 21-step ramp.
The point is structural rather than numeric: by deferring the achromatic constraint to a post-correction, the same trained parameters serve both purposes without forcing a compromise basin that satisfies neither.
Figure˜5 shows neutral ramp uniformity and achromatic chroma leakage. MetricSpace’s neutral correction eliminates chroma leakage entirely ($C<10^{-5}$ on the standard gray ramp); GenSpace (Section˜7) achieves the same property structurally rather than by post-correction. CAM16-UCS and Oklab exhibit chroma leakage on the order of $10^{-7}$ for D65 grays; CAM16-UCS’s leakage rises sharply for non-D65 illuminants in its native chromatic-adaptation chain. Table˜2 summarizes all generation properties for MetricSpace v21.
Figure 5: Left: neutral ramp step uniformity. Right: achromatic chroma leakage. Helmlab’s NC drives $C<10^{-5}$ for grays on the 21-step ramp.
Table 2: Generation quality metrics for MetricSpace v21.
@@
Achromatic max $C$ (NC on) $10^{-5}$
Hue RMS (6 sRGB primaries) 26.4°
Hue max error 53.4°
[…]
Round-trip (random XYZ) $\sim$$\times1$0-13
Munsell CV (3,590 iso-hue pairs) 32.0%
Jacobian det (min, 643 sRGB) 0.10
@@
[…]
He et al. 2022 stress(82 pairs) 35.9
MacAdam stress(128, held out) 19.51
Human Feedback stress(screen) 23.26
The achromatic guarantee (NC enabled) means gray gradients and lightness ramps interpolate without color artifacts, which is necessary for design-token generation and palette export. The hue RMS of 26.4° in v21 is worse than v20b’s 18.1°: v21’s distance-optimal basin pulled the chromatic plane out of geometric alignment with HSL primaries. The rigid rotation (Section˜3, Stage 11) is an isometry of the metric and so its choice is free with respect to STRESS; we keep $\varphi=-28.2$° as the default for backwards compatibility with v20b deployments.
User-tunable display alignment.
Since v0.12.2 the rotation is exposed as a public parameter (display_phi_deg in the JSON checkpoint, or the ab_rotate_deg / abRotateDeg constructor argument in Python and JavaScript respectively). The default of 28.2-28.2° matches the values reported in this paper. A re-measured minimax-optimal angle for v21 is $\varphi^{\ast}\approx-11.75$°, which reduces the maximum primary hue error from 53.453.4° to 36.936.9° (RMS unchanged at 26.426.4°, since rotation cannot redistribute the spread of v21’s basin). We do not change the default to $\varphi^{\ast}$ because (i) the RMS gain is negligible, (ii) the v20b heritage value is already deployed in client code, and (iii) applications that need accurate primary alignment should use GenSpace (Section˜7), where geometry is preserved by construction. The rotation is exposed as a configuration knob so MetricSpace can be tuned to local Lab-axis conventions without retraining. The Jacobian was evaluated on a $64\!\times\!64\!\times\!64$ sRGB grid via finite differences; the minimum determinant (0.10) is positive throughout, so the map is locally invertible and orientation-preserving across the entire sRGB gamut.
@@ 6.4 Ablation Study
Table˜3 isolates the contribution of key components using frozen ablation: each row disables one component while keeping all other parameters at their v21-optimized values. This measures the marginal contribution of each component given the current parameter setting, not its isolated effect; a re-optimised ablation would converge to a different basin and is reported separately for the most impactful stages.
Table 3: Frozen ablation on v21: effect of disabling components without re-optimisation.
Configuration COMBVD stress $\Delta$
Full MetricSpace v21 22.48
Euclidean distance only $\sim$27.6 +5.1+5.1
No H-K embedding $\sim$26.5 +4.0+4.0
No hue correction (Stage 4) $\sim$37.9 +15.4+15.4
No dark-L compression $\sim$22.7 +0.2+0.2
Rotation $\varphi=-28.2$° 22.48 0.00
No rotation ($\varphi=0$) 22.48 0.00
With Euclidean distance only, MetricSpace’s space-component reaches stress $\approx 27.6$, already improving on CIEDE2000 (29.20) by 5.5%. The full metric brings it to 22.48, recovering a further $\sim$5 stress points. Among individual stages, hue correction is the most critical ($\Delta\approx+15.4$); H-K embedding contributes $\sim$+4+4. These large frozen-ablation deltas reflect the co-optimization of all parameters — disabling one stage disrupts the balance — so they should be read as “marginal value of this component conditional on the rest being v21-fitted”, not as “what would this component buy in isolation”. The rotation rows empirically confirm the proof in Section˜3.2: the distance metric is exactly invariant under any rigid $(a,b)$ rotation. This is what allows us to choose $\varphi$ purely on hue alignment without trading off stress.
@@ 6.5 Sub-Dataset Performance
Table˜4 breaks down stress by combvd sub-dataset. MetricSpace v21 outperforms CIEDE2000 on three of six sub-datasets (BFD-P(D65), BFD-P(M), and tied on BFD-P(C)) and trails on three (LEEDS, RIT-DuPont, WITT) where CIEDE2000 was specifically tuned. The overall stress lead (22.48 vs 29.20) comes from the dominant performance on BFD-P(D65)+BFD-P(M), which together account for 2,5762{,}576 of 3,8133{,}813 pairs ($\sim$68%). This is what we would expect from a model whose loss is a 50/50 mix of full and balanced sub-dataset stress: the optimum chases the larger sub-datasets where the data signal is strongest, and pays a small price on small, idiosyncratic sets where CIEDE2000’s hand-engineered tuning already exploits the noise structure.
Table 4: stress on individual combvd sub-datasets (v21 production checkpoint).
Sub-dataset NN v21 CIEDE2000
BFD-P(C) 200 29.08 29.08
BFD-P(D65) 2,028 21.54 24.09
BFD-P(M) 548 21.75 35.23
LEEDS 307 21.84 19.25
RIT-DuPont 312 21.90 19.47
WITT 418 30.93 30.22
6.6 Blue-Band: a v20b artefact, absorbed in v21
A previous version of this work (v20b parameters, paper v2) reported a severe gradient non-uniformity in the blue–cyan region: a red-to-blue gradient interpolated in v20b Lab coordinates showed a $51.4\times$ max/min step-size ratio (when measured with CIEDE2000), compared to $1.3\times$ for the same gradient in CIE Lab. The v20b training procedure used an explicit blue-band penalty (weighted sum of LEEDS, RIT-DuPont and a synthetic blue-segment stress) to suppress this. The resulting fix reduced the ratio to $5.8\times$ (an $8.9\times$ improvement) at a cost of only +0.08+0.08 stress on combvd.
[…]
7 GenSpace: a Generation-Optimized Companion
A color space optimized for distance prediction does not necessarily produce usable coordinates for color generation. MetricSpace’s distance-fit pushes the chroma scaling, hue corrections, and pair weights into shapes that are quantitatively justified but produce visually unbalanced gradients — particularly across the blue-magenta range where the high-chroma chroma_power term and the hue-modulated $S_{C}$ both spike. Adding generation constraints to the MetricSpace loss damaged stress without fully fixing gradients. We instead train a separate, simpler space targeted at generation quality and ship both side-by-side.
Pipeline.
GenSpace v0.11.1 follows a stripped-down version of the MetricSpace shape:
[…]
$\;\to$ $C^{0.978}\;\to\;\mathrm{PW}_{L}^{19}\;\to\;\mathrm{enr}_{h}(L)\;\to\;\mathrm{Lab}$
There is no Helmholtz–Kohlrausch coupling, no hue-Fourier rotation, no neutral-correction LUT, and no rigid post-rotation; the achromatic axis is held at $a=b=0$ structurally via a smooth neutral-blend that activates only when channel spread is sub-microscopic relative to the channel mean.
Why depcubic instead of cube-root.
Cube-root has infinite derivative at zero ($\partial f/\partial x=\tfrac{1}{3}x^{-2/3}\to\infty$), which produces numerically unstable gradients near black and is the root cause of the well-known cubic-cusp gamut artefact in OKLab. The depressed cubic $y^{3}+\alpha y=x$ with $\alpha=0.021$ has finite derivative everywhere, fits cube-root closely for $|x|\gg\alpha$, and admits a closed-form forward (Cardano / Halley refinement) and a polynomial inverse. The hole count in the sRGB gamut at $h\!\approx\!265$° drops from 46 (OKLab) to $5$ (GenSpace), each $\sim 0.001$ chroma wide — effectively sub-pixel.
Why a 19-point dense piecewise-linear $L$ correction.
A 9-point sparse correction (used in v0.10.x) was tuned by least-squares against Munsell value samples. The 19-point dense version is fitted against 128 evenly-spaced lightness levels with a stricter monotonicity constraint; on the 43-test ColorBench gradient suite this produces 4 additional wins versus OKLab and 1 fewer loss compared to the sparse version, with no regressions.
Why $C^{0.978}$.
A chroma power of $0.978$ slightly contracts high-C distances, which on the gradient-CV (coefficient of variation across step sizes) metric reduces the median CV from $\sim$38% to $\sim$37.5% and the worst-case CV from 412% to 378%. The exponent was selected by direct grid search on gradient CV; ColorBench stress on combvd rises noticeably with chroma_power <0.95<0.95, so 0.978 is the strongest contraction that does not degrade distance prediction catastrophically.
Why L-gated hue enrichment.
A blue-to-white interpolation in GenSpace without enrichment exhibits a visible purple shift around $L\!\approx\!0.6$ (a known artefact of the blue-region gamut fold). A localised hue rotation of amplitude $0.058$ centred at h=264.5h=264.5° ($\sigma=0.7$ rad), gated to the lightness window $L\!\in\![0.37,1.0]$ via a $\sin^{2}$ ramp, removes the purple shift. The gate width is the largest that does not introduce hue drift in non-blue gradients.
Result.
[…]
Figure 6: GenSpace v0.11.1 vs OKLab — per-category breakdown across all 90 ColorBench metrics. Headline tally is 65W / 9L / 16T (Table˜5). Categories ordered by total metric count.
Table 5: GenSpace v0.11.1 vs OKLab — head-to-head on ColorBench. Independent = Hung-Berns, Ebner-Fairchild, Pointer-gamut metrics (held out from optimizer). Visual breakdown in Figure˜6.
Category GenSpace OKLab Ties
Gamut geometry 24 0 3
Application 9 0 3
Gradient quality 7 3 1
Independent 6 1 0
Perceptual 5 0 0
Structural 4 2 2
Hue 2 0 0
Achromatic 2 0 0
Advanced 2 0 4
Special 2 1 0
Banding 1 0 1
Accessibility 1 1 0
Numerical stability 0 1 2
Total 65 9 16
When to use which.
MetricSpace is the right answer for accessibility checking, tone matching, color quantization, and any application where “how different are these two colors to a human” is the central question. GenSpace is the right answer for palette ramps, gradient authoring, color-mix in CSS, gamut mapping, and animation timing — anywhere the gradient’s step uniformity matters more than its absolute distance value. The two spaces are convertible (both sit on D65 XYZ); designers can specify in one and measure in the other.
@@ 8 Practical Adoption
Figure˜7 shows the sRGB gamut boundary in MetricSpace coordinates at three lightness levels. The boundaries are smooth and well-behaved. The two Helmlab spaces ship with cross-language implementations and integrations into the major web color toolchains:
@@
Cross-language implementations.
The reference implementation is a Python package (helmlab on PyPI, NumPy/SciPy backend, 308 unit tests) covering both spaces, every transform stage, and the distance metric. A pure-JavaScript port (helmlab on npm, zero runtime dependencies, ESM) ships the same checkpoints with 196 unit tests; the JS port has been validated against the Python implementation to a maximum coordinate drift of $\sim$$\times1$0-14 across 1,0001{,}000 random sRGB samples.
Integration into Color.js.
A pull request adding all four Helmlab spaces — helmlab-metric, helmgen, helmgenlch (cylindrical GenSpace), and deltaEHelmlab as a registered distance method — was merged into Color.js (PR #722, May 2026). Color.js is the reference JS color library used by the W3C CSS working group’s color tools and several browser developer tools; having both spaces accessible via the standard new Color("helmlab-metric", [...]) constructor brings Helmlab into the same toolchain as Oklab, Lab, OKLCh, Jzazbz, and ICTCP. CSS Color Level 4 color(--helmlab-metric ...), color(--helmgen ...), and color(--helmgenlch ...) syntax round-trips through the parser and serialiser.
PostCSS plugin.
postcss-helmlab provides build-time CSS function-form transformation: authors write color: helmlab(0.78 0.52 -0.20) or linear-gradient(in helmlab, red, blue) in source CSS and the plugin emits an rgb() fallback plus @supports-wrapped color(display-p3 ...) and color(rec2020 ...) overrides. The plugin handles helmlab(), helmlch(), helmgen(), and helmgenlch() alongside the standard linear-gradient(in <space>, ...) and color-mix(in <space>, ...) CSS functions. The package ships a dual ESM/CJS build for compatibility with both modern bundlers and Next.js’s PostCSS pipeline.
@@
Binary-search chroma reduction along the L axis maps out-of-gamut Helmlab coordinates to the nearest in-gamut sRGB, Display P3, or Rec. 2020 color, preserving hue and lightness.
@@
Lightness ramps and hue rings at fixed chroma produce perceptually-spaced palettes. Semantic scales (Tailwind-style 50–950) interpolate $L$ between 0.97 and 0.10 with chroma/hue preserved. The default scaling is computed in GenSpace because the gradient-uniformity metrics dominate authoring quality.
@@
The surround parameter $S$ is architecturally present but in v21 is held at zero. The current dark/light mode adaptation uses a soft L inversion across the GenSpace L axis, with contrast enforcement on the resulting pair. A future MetricSpace release with trained $S$ parameters will replace this heuristic.
@@
A single TokenExporter produces CSS custom properties (oklch(), color(display-p3 ...), color(--helmgenlch ...)), Android XML (using HCT tone/chroma/hue semantics for Material Design compatibility), iOS Swift (Display P3), Tailwind config, and raw JSON. Each target reads from the same in-memory token tree, so a token edit propagates to every output format simultaneously.
@@ 9 Limitations and Future Work
    Model complexity. 72 trained parameters is considerably more than Oklab ($\sim$15 designed parameters). Each stage has a perceptual motivation, but the overall model is harder to interpret than simpler spaces. The architecture additionally reserves 22 surround/hue-modulated slots that are held at zero in v21.
@@
[…]
@@
    Hue alignment. RMS hue error of 26.4° is a substantial compromise; the worst primary (cyan) is 53.4° off the HSL reference. v21’s training pushed parameters into the distance-optimal basin at the cost of geometric uniformity — a trade we accept because the distance metric is the headline product, but applications that need accurate hue specification should compose MetricSpace with an explicit hue-rotation post-step or use GenSpace (where the geometry is preserved by construction).
@@
    Surround dependency. The surround parameter $S$ is architecturally present but in v21 the 11 surround-suffix parameters are all held at zero (training data is average-surround only). Current dark/light adaptation uses a heuristic L-inversion in GenSpace with contrast enforcement.
@@
    Sub-dataset variation. v21 trails CIEDE2000 on three of six combvd sub-datasets (LEEDS, RIT-DuPont, WITT). The lead on BFD-P(D65) and BFD-P(M) is large enough to dominate the aggregate, but designers comparing spaces on a single small sub-dataset may pick different winners.
@@
    Not general-purpose. Helmlab is optimized for UI workflows (sRGB/P3 gamut, screen viewing, D65 illuminant). It is not designed for print, photography, illuminant-A scenes, or spectral applications. CAM16 remains the appropriate choice for arbitrary viewing conditions.
@@
  9. 9.
    GenSpace sRGB round-trip. GenSpace’s sRGB round-trip is $\sim 5.6\!\times\!10^{-8}$, six orders of magnitude looser than MetricSpace’s. This is below the visual threshold and matches OKLab’s worst-case round-trip in the same regions, but it is looser than the other GenSpace round-trips (P3, Rec. 2020 are at machine epsilon). The cause is the L-gated enrichment Newton inversion in the blue band; an analytic inverse for the enrichment is open work.
For future work, we plan to train the surround parameters on paired light/dark viewing data, to add an analytic enrichment inverse for GenSpace, and to expand the training set with targeted human feedback using the bidirectional optimisation framework already present in the codebase.